Pure acid is to be added to a 5% acid solution to obtain 95L of 30% solution. What amounts of each should be used?
How many Liters of 100% acid should be used to obtain the solution?

Answer by stanbon(75887)

Pure acid is to be added to a 5% acid solution to obtain 95L of 30% solution. What amounts of each should be used?
----
Equation:
acid + acid = acid
0.05x + 1.0(95-x) = 0.30*95
----
Multiply thr by 100 to get:
5x + 100*95 - 100x = 30*95
---
-95x = -70*95
---
x = 70L (amt of 5% solution needed)
---
95-x = 25L (amt of 100% solution needed)
==================================================
Cheers,
Stan H.
============



How many Liters of 100% acid should be used to obtain the solution?
---------------------